# Safe and Responsible Large Language Model : Can We Balance Bias Reduction and Language Understanding in Large Language Models?

### v3

@@ 3.3 Baselines
Probability-based Baselines: We utilize fine-tuning models such as T5 [32], Flan T5 [33], and BART [34] to compute the probability of content being toxic/harmful on the testsets.
Prompt-based Baselines: This method involves guiding LLMs using specific prompts or prompt tuning. The baselines included in this family of models are Llama2 variants [23], Falcon7B [35], and GPT-2 [36] models. Llama2 and Falcon models are promt-tuned, whereas openAI 666missing GPT-3.5 and GPT-4 are tested using 2-shot prompts.
For the baselines, we use the following terms throughout the paper: Fine-tuning, which involves adjusting pre-trained model weights through additional training on a specific dataset to improve its performance on related tasks; Prompt-tuning, which involvess designing input prompts to prepare a model with fixed weights towards desired outputs [37]; Instruction fine-tuning, which enhances a model ability to follow instructions and respond appropriately by training it with explicit commands [33]; and Prompts with demonstrations, which involves providing input-output example pairs along with prompts to guide the model towards accurate understanding and generation of responses [38].
@@ 3.4 Evaluation Metrics
1.
Probability-based scoring: Employs the Perspective API [39] to measure toxicity probabilities.
2.
LLM-based scoring: Utilizes OpenAI moderation API [40] which provides confidence scores. These scores indicate the likelihood of content being unsafe, and a preset threshold (we used 0.5) is used to determine whether content exceeds acceptable safety limits.
@@
1.
Language Modeling Score (LMS): Measures complete language understanding with a perfect score of 100 indicating full knowledge retention.
2.
Stereotype Score (SS): Assesses the model bias by measuring its tendency toward stereotypical or anti-stereotypical terms. A score of 50 represents a neutral stance, while deviations from 50 indicate a bias toward stereotype or anti-stereotype terms.
3.
Idealized Context Association Test (ICAT): Integrates LMS and SS to simultaneously evaluate language competence and bias neutrality. An ideal ICAT score is 100, which a model would achieve if it scores a perfect LMS of 100 (indicating excellent language understanding) and an SS of 50 (showing no bias towards or against stereotypes).
Content Diversity and Style Metrics Content-Length Entropy Normalization (CLEN): Adapted from the HolisticBias study [41], this metric involves a style classifier from ParlAI [42] that detects attributes like sentiment and writing style. CLEN measures the entropy of sentence lengths to assess stylistic diversity. A higher CLEN score indicates better alignment with desired traits (e.g., kindness), suggesting stylistic consistency in benign contexts.
@@
4 Results
We conduct these experiments to address our research questions: (1) whether we are able to reduce unsafe content generation while preserving the knowledge or language understanding of our model, and (2) whether instruction fine-tuning adds value to the model performance and capability over the base models.
4.1 Evaluating the Effectiveness of SR in Comparison with Various Language Models Across Different Datasets
We evaluated the performance of SR by comparing it against state-of-the-art models using three distinct test sets for toxicity and harmful content reduction. The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text (RQ2).
@@ 4 Results
| Initial score | 57.82 | 68.18 | 68.82 | 69.78 | 59.34 | 65.29 |
| T5 (770M) (FT) | 23.81 | 39.83 | 33.05 | 28.10 | 26.99 | 30.71 |
| BART (406M) (FT) | 21.34 | 27.92 | 24.39 | 27.10 | 22.15 | 28.28 |
| Llama2 (7B) (PT) | 06.05 | 07.18 | 03.88 | 06.10 | 07.04 | 08.92 |
| Llama2 (13B) (PT) | 06.12 | 07.20 | 04.19 | 07.10 | 07.20 | 08.50 |
| Falcon (7B) (PT) | 18.94 | 26.10 | 02.36 | 10.34 | 19.00 | 27.21 |
| GPT-3.5 (prompt) | 08.20 | 10.10 | 27.34 | 29.10 | 09.35 | 11.76 |
| GPT-4 (prompt) | 06.29 | 06.18 | 09.29 | 06.33 | 07.93 | 07.84 |
| SR | 06.01 | 05.92 | 04.40 | 05.10 | 07.36 | 06.89 |
The results in Table 5 reveal that our instruction fine-tuned SR model that is build on the top of already safe Llama2B-chat model, consistently outperformed other models in terms of low toxicity and content moderation scores. The base Llama2chat models with prompt tuning demonstrated strong performance, particularly the 7B and 13B parameter versions, which exhibited significantly lower toxicity percentages compared to others. Model size seems to impact results, as we observe Llama2-7B performs marginally better than Llama2-13B, which align with original Llama2 study also [23]. The LLMs such as Llama2, Falcon and ours performs better than pre-trained smaller langage models like T5 and BART. Meanwhile, GPT-4 showcased impressive results in the prompt-based few-shot learning category.
Main Finding: The results primarily address RQ2, demonstrating the added value of instruction fine-tuning when applied to inherently safer models. The result also shows that LLMs more effectively recognize and handle toxic content compared to smaller models.
4.2 Evaluating the Effectiveness of SR in Comparison with Various Language Models Across Different Demographics
@@
We assessed various models for their effectiveness in mitigating toxicity across different demographics. Methods included fine-tuning smaller language models, prompt-tuning LLMs (Llama2-7B-chat, Falcon7B), and instruction fine-tuning for SR on our CMD dataset. GPT-3.5/4 were used with 2-shot demonstration prompts to produce safe variations. The models were then used to debias the Toxigen test set, with toxicity measured using ToxiGen-RoBERTa [6]. Probability scores for each sentence were averaged and presented as percentages.
The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases across different demographics in text (RQ2).
| Demo. | Original Toxicity | Llama 2-7B | Llama 2-13B | Falcon -7B | T5 | BART | GPT-3.5 | GPT-4 | SR |
| Women | 92.60 | 5.01 | 12.15 | 13.92 | 25.74 | 24.10 | 3.38 | 1.02 | 3.19 |
| Mental Disability | 90.45 | 2.28 | 1.71 | 7.28 | 18.27 | 18.29 | 3.65 | 1.12 | 1.24 |
| LGBTQ | 86.58 | 2.21 | 6.85 | 11.13 | 21.89 | 20.01 | 3.24 | 0.67 | 1.92 |
| Black | 3.08 | 10.21 | 12.21 | 26.35 | 26.01 | 4.12 | 1.66 | 1.04 | |
| Chinese | 90.48 | 2.03 | 3.76 | 8.23 | 17.68 | 16.74 | 3.25 | 1.46 | 0.98 |
| Asian | 99.19 | 1.98 | 3.80 | 6.02 | 16.77 | 15.10 | 4.10 | 1.23 | 1.37 |
| Native American | 98.27 | 2.34 | 5.52 | 11.62 | 20.96 | 19.35 | 3.81 | 1.47 | 1.85 |
| Middle Eastern | 91.54 | 2.89 | 6.11 | 8.34 | 23.47 | 23.45 | 4.06 | 1.01 | 1.93 |
| Muslim | 94.46 | 3.18 | 10.18 | 13.98 | 23.79 | 24.50 | 3.79 | 1.12 | 1.87 |
| Physical Disable | 82.84 | 1.58 | 1.52 | 6.82 | 18.82 | 17.02 | 3.18 | 0.92 | 0.59 |
| Mexican | 87.48 | 6.72 | 17.03 | 14.21 | 34.27 | 33.56 | 3.80 | 1.22 | 4.28 |
| Jewish | 81.96 | 3.78 | 12.54 | 15.53 | 23.28 | 26.39 | 3.78 | 1.19 | 2.78 |
| Latino | 84.84 | 4.15 | 15.42 | 15.87 | 29.45 | 30.12 | 3.57 | 1.34 | 2.24 |
@@
The results in Table 6 show that GPT-4 and SR consistently outperformed other models by generating texts that exhibited the least amount of toxic content across diverse demographic groups. Specifically, GPT-4 exhibited the lowest toxicity percentages for groups such as women (1.02%), LGBTQ (0.67%), and those with physical disabilities (0.92%). SR also demonstrated strong performance, particularly among groups like those with mental disabilities (1.24%) and the Chinese demographic (0.98%). In contrast, T5 and BART were the least effective, producing the highest percentages of unsafe content across most demographic categories.
@@
Main Finding: Instruction fine-tuning models perform better than simple fine-tuning or prompt-tuning, addressing RQ2. Closed-source GPT-4 is a strong baseline in reducing toxicity, while smaller models like T5 and BART struggle in this regard.
4.3 Evaluating the Effectiveness of SR in Comparison with Various Language Models for Stereotypes
@@
We assessed SR alongside other leading models using the StereoSet [26] to determine bias across four demographic categories: gender, profession, race, and religion. Our analysis utilized the Intrasentence task from StereoSet, chosen over the Intersentence Test due to its suitability for detailed language generation and bias evaluation. We employ well-established baselines such as Flan-T5, GPT-2 Large, and DialogGPT, each fine-tuned on our dataset. These models were selected based on their adaptability and performance on similar tasks.
The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text (RQ2) while retaining language understanding (RQ1).
@@
| Model | LMS | SS | ICAT | LMS | SS | ICAT |
| Flan-T5base (FT) | 87.84 | 56.70 | 76.07 | 89.01 | 59.64 | 71.85 |
| Flan-T5medium (FT) | 88.63 | 55.31 | 79.22 | 84.32 | 61.79 | 64.44 |
| Flan-T5large (FT) | 92.55 | 65.25 | 64.32 | 91.36 | 61.62 | 70.13 |
| GPT2large (FT) | 80.77 | 70.93 | 46.96 | 79.99 | 64.34 | 57.05 |
| DialoGPTlarge (FT) | 82.50 | 61.29 | 63.87 | 79.87 | 58.72 | 65.94 |
| Llama2 7B (PT) | 92.64 | 65.30 | 64.29 | 91.30 | 63.31 | 67.00 |
| Llama2 13B (PT) | 91.20 | 66.44 | 61.21 | 90.25 | 64.22 | 64.58 |
| SR (IFT) | 90.05 | 58.47 | 74.80 | 90.58 | 62.25 | 68.39 |
@@
| Model | LMS | SS | ICAT | LMS | SS | ICAT |
| Flan-T5base (FT) | 86.38 | 68.23 | 54.89 | 83.54 | 69.70 | 50.63 |
| Flan-T5medium (FT) | 83.47 | 62.52 | 62.57 | 83.54 | 62.12 | 63.29 |
| Flan-T5large (FT) | 91.48 | 62.16 | 69.23 | 96.20 | 80.26 | 37.98 |
| GPT2large (FT) | 69.43 | 68.35 | 43.95 | 66.09 | 75. | 31.76 |
| DialoGPTlarge (FT) | 83.44 | 60.51 | 65.90 | 84.91 | 67.23 | 55.65 |
| Llama27B (PT) | 92.27 | 65.01 | 64.57 | 92.10 | 61.05 | 71.75 |
| Llama2 13B (PT) | 91.28 | 66.78 | 60.65 | 91.23 | 62.92 | 69.48 |
| SR (IFT) | 92.44 | 61.76 | 70.70 | 93.68 | 61.35 | 72.41 |
The results presented in Table 7 demonstrate the superior performance of SR in reducing stereotypical biases across multiple dimensions: Gender, Profession, Race, and Religion. In particular, SR achieves an impressive balance between diminishing biases and preserving language modeling capabilities, evidenced by higher ICAT scores across various demographics. SR exhibits balanced performance in the Gender and Profession categories, with high Language Modeling Scores (LMS) of 90.05 and 90.58 respectively, while maintaining low stereotype scores (SS) and high ICAT scores. In the Race and Religion categories, SR not only achieves high LMS scores (92.44 for Race and 93.68 for Religion) but also records the lowest SS, underlining its effectiveness in managing sensitive content with minimal bias
@@
4.4 Evaluation of SR on Text Style Factors

### v5

@@ 3.3 Baselines
Encoder-Decoder Baselines: We utilize fine-tuning on smaller encoder-decoder language models such as T5 [40], Flan T5 [33], and BART [41]. We used their large versions, which are still considered smaller than LLMs. These models are chosen for their capabilities in generating coherent and contextually relevant text. We fine-tuned these models’ weights based on our unsafe-safe pair content data, the idea is to enable these models to learn the nuances and patterns associated with biased content, for the generation of safe and bias-free text.
Prompt-based Baselines: We use state-of-the-art LLMs, such as Llama2-chat variants [26], Llama3.1, Llama 3.2, Falcon7B [42], GPT-2 [43], and OpenAI’s GPT-3.5 and GPT-4 models 777https://openai.com/ using prompt-based techniques. This approach involves guiding the LLMs with specific prompts to generate desired outputs. We selected few-shot learning for these experiments, based on our preliminary analysis that demonstrated better performance compared to zero-shot learning. We opted for 2-shot learning to balance computational costs and performance, especially considering the cost implications of using paid models like GPT-3.5 and GPT-4. We also instruction fine-tuned Llama2,-7B-chat, Llama3-8B-instruct, Llama3.1-8B-instruct and Llama3.2-1B-instruct in this work to show SR working.
@@ 3.4 Evaluation metrics
Probability-based scoring: We employ the Perspective API [44] to measure the probability that a comment is perceived as toxic. This model provides a score indicating the likelihood that a text will be considered toxic or non-toxic. We specified the default threshold value.
LLM-based scoring: We utilizes OpenAI moderation API [45] which provides confidence scores. These scores indicate the likelihood of content being unsafe, and a preset threshold (we used 0.5) is used to determine whether content exceeds acceptable safety limits.
We also used the knowledge retention metric from DeepEval [46] that determines whether a LLM is able to retain factual information presented throughout a conversation. It takes the original text (unsafe example) and its variation (safe version) to determine whether the output response indicates an inability to recall said knowledge.
We used GPT-4 as the backend LLM to evaluate LLM-based scoring. For our problem, a lower () toxicity and content moderation scores are considered good, indicating reduced toxicity and improved content quality.
@@
Language Modeling Score (LMS): LMS measures complete language understanding with a perfect score of 100 indicating full knowledge retention. A higher () score is considered better.
Stereotype Score (SS): SS assesses the model bias by measuring its tendency toward stereotypical or anti-stereotypical terms. A score of 50 represents a neutral stance, while deviations from 50 indicate a bias toward stereotype or anti-stereotype terms.
Idealized Context Association Test (ICAT): ICAT integrates LMS and SS to simultaneously evaluate language competence and bias neutrality. A higher () score is considered better. An ideal ICAT score is 100, which a model would achieve if it scores a perfect LMS of 100 (indicating excellent language understanding) and an SS of 50 (showing no bias towards or against stereotypes).
Content Diversity and Style Metrics Content-Length Entropy Normalization (CLEN): CLEN metric is adapted from the HolisticBias study [47], this metric involves a style classifier from ParlAI [48] that detects attributes like sentiment and writing style. CLEN measures the entropy of sentence lengths to assess stylistic diversity. A higher () CLEN score indicates better alignment with desired traits (e.g., kindness), suggesting stylistic consistency in benign contexts.
@@
4 Results and Discussion
We conduct experiments to address three key research questions: (1) Can we reduce unsafe content generation while retaining the model’s knowledge and language understanding? (2) Is instruction fine-tuning more effective than prompting alone? (3) Does instruction fine-tuning on safety-tuned models further enhance their performance compared to base models?
4.1 Performance evaluation across different datasets
We evaluated the performance of SR by comparing it against state-of-the-art models using three distinct test sets for toxicity and harmful content reduction. The goal of this experiment is to see if task-specific instruction fine-tuning can reduce biases in language generation.
@@ 4.1 Performance evaluation across different datasets
| Pre-safety Scores | ||||||
| Original Texts | 57.82 | 68.18 | 68.82 | 69.78 | 59.34 | 65.29 |
| Fine-tuned (FT) | ||||||
| T5 (770M) | 23.81 | 39.83 | 33.05 | 28.10 | 26.99 | 30.71 |
| BART (406M) | 21.34 | 27.92 | 24.39 | 27.10 | 22.15 | 28.28 |
| Prompt-based (P) (2-shots setting) | ||||||
| Falcon (7B) | 18.94 | 26.10 | 02.36 | 10.34 | 19.00 | 27.21 |
| GPT-3.5 | 08.20 | 10.10 | 27.34 | 29.10 | 09.35 | 11.76 |
| GPT-4 | 06.29 | 06.18 | 09.29 | 06.33 | 07.93 | 07.84 |
| Llama2 (7B) | 13.05 | 17.18 | 14.88 | 16.10 | 17.04 | 16.92 |
| Llama3 (8B) | 05.92 | 05.85 | 08.95 | 07.88 | 07.20 | 06.95 |
| Llama3.1 (8B) | 05.75 | 05.80 | 08.10 | 07.55 | 06.90 | 06.75 |
| Llama3.2 (1B) | 05.60 | 05.50 | 07.85 | 07.30 | 06.50 | 06.40 |
| Instruction Fine-tuned (IFT) | ||||||
| SR Llama2 | 06.01 | 05.92 | 04.40 | 05.10 | 07.36 | 06.89 |
| SR Llama3 | 05.90 | 05.80 | 04.00 | 04.95 | 07.10 | 06.70 |
| SR Llama3.1 | 05.75 | 05.75 | 03.90 | 04.80 | 06.95 | 06.50 |
| SR Llama3.2 | 05.50 | 05.40 | 03.75 | 04.60 | 06.80 | 06.40 |
Pre-Safety Scores: The results in Table 4 highlight the initial scores on the original texts. These scores, obtained using the Perspective API and OpenAI content moderation tools, indicate high levels of toxicity and biased content in the raw text, emphasizing the need for interventions to ensure safer text generation.
Post-Safety Scores: After applying safety techniques—including fine-tuning, prompting, and instruction fine-tuning—there is a marked reduction in toxicity and harmful content. Our instruction fine-tuned SR model, built on top of the safety-optimized Llama models, consistently achieves the lowest toxicity and content moderation scores across all test sets. Notably, decoder-only models (Llama, Falcon, GPT) outperform encoder-decoder models (T5, BART) in terms of toxicity reduction.
The base Llama2/3/3.1/3.2 models with prompt-based methods demonstrate strong performance, achieving lower toxicity and bias levels compared to other baseline models. Similarly, GPT-4 exhibits impressive results in the prompt-based few-shot learning category, reaffirming the utility of large-scale decoder-only architectures in safe text generation.
Main Findings: The results directly address RQ2 by demonstrating the efficacy of instruction fine-tuning in enhancing safety and reducing toxicity, as evidenced by the performance of the SR model. Furthermore, they substantiate RQ3 by showing that instruction fine-tuning applied to inherently safer models, such as the Llama series, leads to significant performance improvements. Additionally, the analysis underscores that autoregressive, decoder-only LLMs are more effective in recognizing and mitigating toxic content compared to smaller encoder-decoder models like T5 and BART.
In our later experiments, we use SR with Llama3.2. This decision is based on the marginal differences observed among the Llama variants, with Llama3.2 demonstrating greater stability and a slight performance edge.
4.2 Performance evaluation across different demographics
@@
We evaluated various models to mitigate toxicity across different demographics. Our methods included fine-tuning smaller language models like T5 and BART, LLMs such as Llama, Falcon7B, and GPT-3.5/4, using prompts ,and instruction fine-tuning for SR on our CMD dataset. The models were tested on the Toxigen test set, with toxicity levels assessed using ToxiGen-RoBERTa [6], a model specifically trained for this dataset. We presented the toxicity scores as averaged probabilities in percentage form.
| Demographic Group | Original Toxicity | T5 | BART | Llama 3.2 -1B | Falcon-7B | GPT-3.5 | GPT-4 | SR (IFT) |
| Women | 92.60 | 25.74 | 24.10 | 4.98 | 13.92 | 3.38 | 1.02 | 3.08 |
| Mental Dis. | 90.45 | 18.27 | 18.29 | 2.25 | 7.28 | 3.65 | 1.12 | 1.12 |
| LGBTQ | 86.58 | 21.89 | 20.01 | 2.34 | 11.13 | 3.24 | 0.67 | 1.78 |
| Black | 90.48 | 26.35 | 26.01 | 2.98 | 12.21 | 4.12 | 1.66 | 1.02 |
| Chinese | 86.52 | 17.68 | 16.74 | 1.99 | 8.23 | 3.25 | 1.46 | 0.75 |
| Asian | 99.19 | 16.77 | 15.10 | 1.73 | 6.02 | 4.10 | 1.23 | 1.29 |
| Native Amer. | 98.27 | 20.96 | 19.35 | 2.12 | 11.62 | 3.81 | 1.47 | 1.83 |
| Middle Est. | 91.54 | 23.47 | 23.45 | 2.18 | 8.34 | 4.06 | 1.01 | 1.83 |
| Muslim | 94.46 | 23.79 | 24.50 | 2.98 | 13.98 | 3.79 | 1.12 | 1.58 |
| Physical Dis. | 82.84 | 18.82 | 17.02 | 1.56 | 6.82 | 3.18 | 0.59 | 1.08 |
| Mexican | 87.48 | 34.27 | 33.56 | 4.01 | 14.21 | 3.80 | 1.22 | 1.20 |
| Jewish | 81.96 | 23.28 | 26.39 | 3.10 | 15.53 | 3.78 | 1.19 | 2.26 |
| Latino | 84.84 | 29.45 | 30.12 | 4.03 | 15.87 | 3.57 | 1.34 | 2.10 |
@@
The results presented in Table 5 indicate that GPT-4, and SR in both variations (prompting and IFT) consistently outperformed other models in generating texts with minimal toxic content across diverse demographic groups. Specifically, GPT-4 recorded the lowest toxicity percentages for groups including women (1.02%), LGBTQ (0.67%), Physical disability (0.59%)and other demographics, as highlighted in Table 5. SR also showed robust performance, especially among groups such as individuals with the Chinese, Mexicans and Black demographics. In contrast, models such as T5 and BART were less effective, exhibiting the highest percentages of unsafe content across most demographic categories.
@@
Main Finding: Closed-source GPT-4 is a strong model in reducing toxicity, while smaller models like T5 and BART struggle in this regard. Instruction fine-tuning models on the top of inherently safe models (e.g. our model SR compared to base model) perform better than simple fine-tuning or prompt-tuning, addressing RQ3.
4.3 Performance evaluation for stereotypes
@@
We assessed SR alongside other leading models using the StereoSet [30] to determine bias across four demographic categories: gender, profession, race, and religion. Our analysis utilized the Intrasentence task from StereoSet, chosen over the Intersentence Test due to its suitability for detailed language generation and bias evaluation. We employ well-established baselines such as Flan-T5, GPT-2 Large, and DialogGPT, each fine-tuned on our dataset. These models were selected based on their adaptability and performance on similar tasks. The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text, while retaining language understanding.
@@
| Model | LMS () | SS | ICAT () | LMS () | SS | ICAT () |
| Flan-T5base FT | 87.84 | 56.70 | 76.07 | 89.01 | 59.64 | 71.85 |
| Flan-T5medium FT | 88.63 | 55.31 | 79.22 | 84.32 | 61.79 | 64.44 |
| Flan-T5large FT | 92.55 | 65.25 | 64.32 | 91.36 | 61.62 | 70.13 |
| GPT2large FT | 80.77 | 70.93 | 46.96 | 79.99 | 64.34 | 57.05 |
| DialoGPTlarge FT | 82.50 | 61.29 | 63.87 | 79.87 | 58.72 | 65.94 |
| Llama3.2 1B (P) | 90.12 | 59.18 | 61.29 | 90.25 | 56.19 | 68.58 |
| SR (IFT) | 91.05 | 54.47 | 76.98 | 91.98 | 62.10 | 74.12 |
@@
| Model | LMS () | SS | ICAT () | LMS () | SS | ICAT () |
| Flan-T5base FT | 86.38 | 68.23 | 54.89 | 83.54 | 69.70 | 50.63 |
| Flan-T5medium FT | 83.47 | 62.52 | 62.57 | 83.54 | 62.12 | 63.29 |
| Flan-T5large FT | 91.48 | 62.16 | 69.23 | 96.20 | 80.26 | 37.98 |
| GPT2large FT | 69.43 | 68.35 | 43.95 | 66.09 | 75. | 31.76 |
| DialoGPTlarge FT | 83.44 | 60.51 | 65.90 | 84.91 | 67.23 | 55.65 |
| Llama3.2 1B (P) | 91.28 | 66.78 | 60.65 | 91.23 | 62.92 | 69.48 |
| SR (IFT) | 93.48 | 52.76 | 74.13 | 93.99 | 62.23 | 77.24 |
The results presented in Table 6 demonstrate the superior performance of SR in reducing stereotypical biases across multiple dimensions: Gender, Profession, Race, and Religion. In particular, SR achieves an impressive balance between reducing biases and preserving language modeling capabilities, which is evidenced by higher ICAT scores across various demographics. SR exhibits balanced performance in the Gender and Profession categories, with high LLMS of 91.05 and 91.98 percentages respectively, while maintaining low SS (close to 50) and high ICAT scores (above 50). In the Race and Religion categories, SR not only achieves high LMS scores (93.48 for Race and 93.99 for Religion) but also records the lowest SS (closer to 50).
@@
4.4 Performance evaluation on text style factors